OpenStudy (anonymous):

Solve, using the substitution method. y – 3x = 7 21 + 9x = 3y

OpenStudy (anonymous):

A. There is no solution. B. There are an infinite number of solutions. C. The solution is (1, 10) D. The solution is (28, 0)

OpenStudy (luigi0210):

Move the 3x: y=3x+7 Now sub in: \[\LARGE 21+9x=3(3x+7)\] Distribute: \[\LARGE 21+9x=9x+21\] Simplify: \[\LARGE 21=21\] Which is a true statement so it infinitely many solutions

OpenStudy (anonymous):

thank you, so B?

OpenStudy (luigi0210):

Yup!

OpenStudy (anonymous):

Thanks!

OpenStudy (luigi0210):

You are very welcome